OpenStudy (anonymous):

What is the vertex of the graph of y = −3x^2 − 4x + 2

OpenStudy (anonymous):

vertex:x = -b/2a ax^2 + bx + c b= -4 a= -3 vertex: x= -(-4)/2(-3) = 4/-6 = -2/3 y = -3(-2/3)2 - 4(-2/3) +2 http://hotmath.com/hotmath_help/topics/vertex-of-a-parabola.html

OpenStudy (anonymous):

* I meant for that last line: y = -3(-2/3)^2 - 4(-2/3) +2

OpenStudy (anonymous):

So y=-2

OpenStudy (anonymous):

tyvm :D